1. Anna’s demand for peaches is given by PA = 200 – 3QA, where QA is the quantity (kilograms) demanded at price PA ($/kilo). Basil’s demand is given by PB = 120 – 2QA, where QB is the quantity (kilograms) demanded at price PB ($/kilo). Each of the two has $2,000 that they can spend if they want to buy something.

a. Suppose the endowments are as follows: Anna has 110 kilograms of peaches, Basil has none. When they trade, who will sell to whom? How many kilos will be sold? At what price?

b. Suppose the endowments are as follows: Anna has no peaches, Basil has 110 kilos. When they trade, who will sell to whom? How many kilos will be sold? At what price?

c. Suppose the endowments are as follows: Anna has 60 kilograms of peaches, Basil has 50 kilos. When they trade, who will sell to whom? How many kilos will be sold? At what price?

d. Summarize, in your own words, the answers to parts a) through
1
Expert's answer
2019-05-31T09:02:53-0400

a. Anna has 110 kilograms of peaches, Basil has none. When they trade, Anna will sell to Basil.

PA = PB, QB = 110 - QA,

200 – 3QA = 120 – 2(110 - QA),

5QA = 300,

QA = 60 kilo, QB = 50 kilos will be sold.

P = 200 - 3*60 = 20.

b. Anna has no peaches, Basil has 110 kilos. When they trade, Basil will sell to Anna.

200 – 3(110 - QB) = 120 – 2QB,

QB = 50 kilo, QA = 60 kilos will be sold.

P = 200 - 3*60 = 20.

c. Suppose the endowments are as follows: Anna has 60 kilograms of peaches, Basil has 50 kilos. There will be no trade.

d. If we summarize the answers, then we can conclude, that in any case of trade the resulting quantities demanded and prices are the same.